# Towards a theory of natural directed paths

### v3

@@ Abstract.
2020 Mathematics Subject Classification:
55U35,68Q85
1. Introduction
Precubical sets are de facto the standard geometric model for directed homotopy for concurrency [2]. In fact, most of them are even non-positively curved in the sense of [12, Definition 1.28 and Proposition 1.29], or in the worst case proper in the sense of [19, page 499]. The motivation for introducing symmetric transverse sets in [6] is to formalize the parallel product with synchronization for process algebra using the associated coskeleton functor [6, Theorem 4.1.8]. Indeed, it is impossible to use the coskeleton functor associated with the category of precubical sets because of its pathological behavior (see [5, Proposition 3.15] and [6, Definition 3.1.3]). However, precubical sets still remain sufficient to model this parallel product by tweaking the coskeleton functor of this category (see [5, Section 3.3]).
Symmetric transverse sets share with precubical sets similar metric and homotopical properties by [11]. Indeed, their geometric realization carries a Lawvere metric structure which enables us to extend Raussen’s notion of (tame) natural $d$-path originally defined for precubical sets [17, Definition 2.14] [19, Definition 5.3] [20, Section 2.9]. Moreover, the full subcategory of representable objects of this presheaf category is c-Reedy in the sense of [18, Definition 8.25], like the box category (see Definition 2.3) for the precubical sets (the box category is even direct Reedy in the sense of [14, Definition 15.1.2]). This makes possible to compare in [11, Theorem 7.4] the natural realization of a symmetric transverse set with other realization functors and to generalize homotopical results proved in [4] and [9] for precubical sets.
The technical contribution of this note is threefold. Firstly, we explain why precubical sets and symmetric transverse sets belong to a larger family of presheaf categories on a thick category of cubes (see Definition 2.10). This family of presheaf categories contains also the symmetric precubical sets of [13] and a new category of non-symmetric transverse sets. Symmetric transverse sets are presheaves on a thick category of cubes which turns out to be, for tautological reasons, the greatest one for the inclusion.
@@ 1. Introduction
Secondly, we prove that all results of [11] are valid for all presheaf categories on a thick category of cubes.
@@
(Section 3 and more specifically Theorem 3.14) All metric and homotopical results of [11] are valid for the category ${\mathcal{A}}^{op}{{\rm{\mathbf{Set}}}}$ of $\mathcal{A}$-sets when $\mathcal{A}$ is a thick category of cubes.
Thirdly, we obtain a statement which coincides with (a part of) [20, Theorem 7.5] when $\mathcal{A}$ is the box category used to define the precubical sets:
@@
(Section 5) The formalization of the parallel composition with synchronisation of [6] using the coskeleton functor of the category of symmetric transverse sets has a category of cube chains which gives the correct space of tame natural $d$-paths up to homotopy.
The formal setting of presheaf category on a thick category of cubes is a first step towards an axiomatization of the notion of tame natural $d$-path. The next step would be to find a way of taking into account the globular version of this notion as it is introduced in [10]. This could lead to a general framework unifying all geometric approaches of directed homotopy for concurrency  [^1] The expression “directed homotopy” has several quite distinct meanings. It is the reason why I add “for concurrency” on purpose..
2. Thick category of cubes
2.1 Definition.
A small category enriched over $([0,\infty],\geqslant,+,0)$ [15] is called a Lawvere metric space. The category of Lawvere metric spaces is denoted by $\mathbf{LvMet}$.
A Lawvere metric space $(X,d)$ is a set $X$ equipped with a map $d:X\times X\to[0,\infty]$ called a Lawvere metric such that:
@@ 2. Thick category of cubes
A map $f:(X,d)\to(Y,d)$ of Lawvere metric spaces is a set map $f:X\to Y$ which is non-expansive, i.e. $\forall(x,y)\in X\times X,d(f(x),f(y))\leqslant d(x,y)$.
2.2 Notation.
@@
Let $[0]=\{()\}$ and $[n]=\{0<1\}^{n}$ for $n\geqslant 1$ equipped with the product order. Let $0_{n}=(0,\dots,0)$ ($n$ times) and $1_{n}=(1,\dots,1)$ ($n$ times) with $n\geqslant 0$. By convention, one has $\{0<1\}^{0}=[0]=\{()\}$. In the sequel, for all $n\geqslant 1$, both the sets $[n]$ and $[0,1]^{n}$ are equipped with the product order. By convention, $[0,1]^{0}$ is a singleton.
@@
Let $x=(x_{1},\dots,x_{n})$ and $x^{\prime}=(x^{\prime}_{1},\dots,x^{\prime}_{n})$ be two elements of $[0,1]^{n}$ with $n\geqslant 1$. Let $\overrightarrow{d}\!_{1}:[0,1]^{n}\times[0,1]^{n}\to[0,\infty]$ be the set map defined by
@@
Let $n\geqslant 0$. The set map $\overrightarrow{d}\!_{1}:[0,1]^{n}\times[0,1]^{n}\to[0,\infty]$ is a Lawvere metric by [11, Proposition 1.5]. It restricts to a Lawvere metric on $\{0<1\}^{n}$.
2.4 Definition.
[6, Definition 2.1.5] A map $f:[m]\rightarrow[n]$ of ${{\rm{\mathbf{PoSet}}}^{+}}$ is cotransverse if
@@
Denote by ^$\widehat{\square}_{S}$ the subcategory of ${{\rm{\mathbf{PoSet}}}^{+}}$ generated by the cotransverse maps.
A cotransverse degeneracy map is a cotransverse map $[n]\to[n]$ for $n\geqslant 2$ which is not one-to-one. Proposition 2.5 and Proposition 2.6 are important for the sequel.
2.5 Proposition.
[6, Proposition 3.1.14] Let $0\leqslant m\leqslant n$. Every cotransverse map $f:[m]\rightarrow[n]$ factors uniquely as a composite $[m]\stackrel{{\scriptstyle\psi}}{{\longrightarrow}}[m]\stackrel{{\scriptstyle\phi}}{{\longrightarrow}}[n]$ with ϕitalic-$\phi\in\square$ and $\psi$ cotransverse.
2.6 Proposition.
Let $f:[m]\to[n]$ be a cotransverse map. Let $\delta:[n]\to[p]$ be a map of normal-$\square$. Suppose that $\delta{f}\in\square$. Then $f\in\square$.
@@
There exists a set map $s:[p]\to[n]$ obtained by removing some coordinates such that $s\delta=\operatorname{Id}_{[n]}$. We deduce that $f=s(\delta{f})$. From $\delta{f}\in\square$, we then deduce that $f\in\square$. ∎
Let $\sigma_{i}:[n]\rightarrow[n]$ be the set map defined for $11\leqslant i\leqslant n-$ 1 and $n\geqslant 2$ by
@@
These maps are called the symmetry maps [13]. The symmetry maps are clearly cotransverse. This is the reason of the $S$ in the notation ^$\widehat{\square}_{S}$.
2.7 Notation.
@@
2.8 Definition.
[6, Definition 2.1.7 and Definition 2.1.12] A category of cubes is a small category $\mathcal{A}$ satisfying the inclusions
@@
The category $\square$, $\square_{S}$ and ^$\widehat{\square}_{S}$ are examples of categories of cubes. The $\square$-sets are the precubical sets [1]. The $\square_{S}$-sets are the symmetric precubical sets (see [13]). The ^$\widehat{\square}_{S}$-sets are the symmetric transverse sets introduced in [11].
@@
By [6, Proposition 2.1.15], the identity map $\operatorname{Id}_{\mathcal{L}_{\mathcal{A}}(K)}$ induces for all precubical sets $K$ a natural inclusion of precubical sets $K\subset\omega_{\mathcal{A}}\mathcal{L}_{\mathcal{A}}(K)$. Let $f,g:K\to L$ be two maps of precubical sets such that $\mathcal{L}_{\mathcal{A}}(f)=\mathcal{L}_{\mathcal{A}}(g)$. Then $\omega_{\mathcal{A}}\mathcal{L}_{\mathcal{A}}(f)=\omega_{\mathcal{A}}\mathcal{L}_{\mathcal{A}}(g)$. Thus
@@
Let $K$ be an object of $\mathcal{A}^{op}{{\rm{\mathbf{Set}}}}$. The set $K([n])$ is denoted by $K_{n}$. An element $x\in K_{n}$ corresponds by the Yoneda lemma to a map of $\mathcal{A}$-sets $x:\mathcal{A}[n]\to K$. The vertex of $x(0_{n})\in K_{0}$ is called the initial state of the $n$-cube $c$ and the vertex $x(1_{n})\in K_{0}$ is called the final state of the $n$-cube $c$. For any map $k:[m]\rightarrow[n]$ of $\mathcal{A}$ and any $\mathcal{A}$-set $K$, denote by $k^{*}:K_{n}\rightarrow K_{m}$ the set map induced by $k$. Let $p\geqslant 0$. The $p$-cube $\mathcal{A}[p]$ is by definition the presheaf $\mathcal{A}(-,[p])$. For any $\mathcal{A}$-set $K$, the data
@@
$\partial\mathcal{A}[n]=\mathcal{A}_{\leqslant n-1}$
@@
The category of cubes ^$\widehat{\square}_{S}$ is thick for tautological reasons. The terminology must be understood as follows. A thick category of cubes $\mathcal{A}$ is morally a thick subcategory of the category of cubes ^$\widehat{\square}_{S}$: it is an analogy with the notion of thick subcategory of a triangulated category. Theorem 2.16 provides other examples of thick categories of cubes.
2.11 Proposition.
@@
${[p]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{k}$$\scriptstyle{f}$${[p]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\delta\in\square}$${[n]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${[n]}$
@@
${[p]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$${[p]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\delta^{\prime\prime}\in\square}$${[p]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{k}$$\scriptstyle{f^{\prime}}$${[q]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\delta^{\prime}\in\square}$${[n]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${[n]}$
Since $\mathcal{A}$ is thick, write $f^{\prime}=\delta^{\prime\prime}g$ with $\delta^{\prime\prime}\in\square$. We obtain $\delta{f}=k=\delta^{\prime}f^{\prime}=\delta^{\prime}\delta^{\prime\prime}g$. By uniqueness of the factorization of Proposition 2.5, we obtain $\delta=\delta^{\prime}\delta^{\prime\prime}$ and $f=g$. We have obtain the map of $(k{\downarrow}j^{n}_{\mathcal{A}})$
${[p]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{k}$$\scriptstyle{f}$${[p]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\delta^{\prime\prime}}$$\scriptstyle{\delta}$${[p]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}}$$\scriptstyle{k}$${[q]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\delta^{\prime}}$${[n]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${[n]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${[n]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${[n]}$
@@
The first statement is [6, Proposition 2.1.14]. Let $n\geqslant 0$. Since $\mathcal{L}_{\mathcal{A}}$ is colimit-preserving, there is a natural map of $\mathcal{A}$-sets ($\square_{<n}$ and $\mathcal{A}_{<n}$ are the full subcategory of $\square$ and $\mathcal{A}$ respectively containing only $[0],\dots,[n-1]$):
@@
The above arrow is an isomorphism by Proposition 2.11 and [16, Theorem 1 p. 213]. ∎
2.13 Proposition.
@@
^=$\widehat{\square}=\{\phi:[m]\to[n]\in\widehat{\square}_{S}\mid\forall\delta:[p]\to[m]\in\square,\phi\delta\hbox{ one-to-one }\Rightarrow\phi\delta\in\square\}$
is closed under composition and contains all identities of ^$\widehat{\square}_{S}$. There are the inclusions ^^$\square\subset\widehat{\square}\subset\widehat{\square}_{S}$. In other terms, the set of maps $\widehat{\square}$ yields a well-defined category of cubes. The only one-to-one set maps of $\widehat{\square}$ are the maps of normal-$\square$.
@@
Let $\phi_{1},\phi_{2}\in\widehat{\square}$ such that $\phi_{1}\phi_{2}$ exists. Let $\delta\in\square$ such that $\phi_{1}\phi_{2}\delta$ exists and is one-to-one. Then $\phi_{2}\delta$ is a one-to-one set map. Thus $\phi_{2}\delta\in\square$, $\phi_{2}$ belonging to $\widehat{\square}$. We deduce that $\phi_{1}\phi_{2}\delta=\phi_{1}(\phi_{2}\delta)\in\square$ since $\phi_{1}\in\widehat{\square}$. This means that $\phi_{1}\phi_{2}\in\widehat{\square}$. For all $\phi=\operatorname{Id}$, one has $\phi\delta=\delta\in\square$. Hence $\widehat{\square}$ contains all identity maps. Finally suppose that $f:[m]\to[n]\in\widehat{\square}$ is one-to-one. Then $f\operatorname{Id}_{[m]}$ is one-to-one, which implies that $f\in\square$. ∎
2.14 Definition.
The $\widehat{\square}$-sets are called transverse sets.
The following maps, introduced in [6, Definition 3.1.11], are examples of cotransverse degeneracy maps. Let $\gamma_{i}:[n]\rightarrow[n]$ be the set map defined for $11\leqslant i\leqslant n-$ 1 and $n\geqslant 2$ by
@@
2.15 Notation.
[6, Theorem 3.1.16] The category of cubes generated by the $\delta_{i}^{\alpha}$, $\sigma_{i}$ and $\gamma_{i}$ is denoted by $\overline{\square}$.
Unlike $\widehat{\square}$, the category of cubes $\overline{\square}$ has a conjectural presentation by generators and relations [6, Proposition 3.1.20 and Conjecture 3.1.21].
2.16 Theorem.
The categories of cubes normal-$\square$, $\square_{S}$, $\widehat{\square}$ and $\overline{\square}$ are thick.
@@
Every map $f:[m]\to[n]$ of $\square$ factors uniquely in ^$\widehat{\square}_{S}$ as a composite $[m]\to[m]\to[n]$ such that the right-hand map $[m]\to[n]$ belongs to $\square$. Since $f=f\operatorname{Id}_{[m]}$, we deduce by uniqueness that the left-hand map belongs to $\square$. Hence the category $\square$ is thick.
Every map $f:[m]\to[n]$ of $\square_{S}$ factors uniquely in ^$\widehat{\square}_{S}$ as a composite $[m]\to[m]\to[n]$ such that the right-hand map $[m]\to[n]$ belongs to $\square$. Since all maps of $\square_{S}$ are one-to-one, the left-hand map $[m]\to[m]$ is one-to-one, hence bijective for cardinality reason. We deduce that the left-hand map is a composite of symmetry maps, which means that it belongs to $\square_{S}$. Thus the category $\square_{S}$ is thick.
Consider a map $f:[m]\to[n]$ of $\widehat{\square}$. It factors in ^$\widehat{\square}_{S}$ as a composite $[m]\to[m]\to[n]$ such that the right-hand map $\delta:[m]\to[n]$ belongs to the box category. Denote by $g:[m]\to[m]$ the left-hand map. Let $\delta^{\prime}:[p]\to[m]\in\square$ such that $g\delta^{\prime}$ exists and is one-to-one. Then $f\delta^{\prime}=\delta(g\delta^{\prime})$ is one-to-one, being the composite of two one-to-one set maps. Since $f\in\widehat{\square}$, we deduce that $\delta{g}\delta^{\prime}\in\square$. Using Proposition 2.6, we deduce that $g\delta^{\prime}\in\square$. We have proved that $g\in\widehat{\square}$. This means that the category of cubes $\widehat{\square}$ is thick.
@@
2.17 Theorem.
Let $\mathcal{A}$ be a thick category of cubes which does not contain the symmetry maps. Then $\mathcal{A}\subset\widehat{\square}$. In other terms, $\widehat{\square}$ is the greatest thick category of cubes for the inclusion which does not contain the symmetry maps.

### v4

# 0000-0003-0287-6252

@@ Abstract
1 Introduction
Presentation
Precubical sets are de facto the standard geometric model for directed homotopy for concurrency DAT_book . In fact, most of them are even non-positively curved in the sense of (zbMATH07226006, , Definition 1.28 and Proposition 1.29), or in the worst case proper in the sense of (MR3722069, , page 499). The motivation for introducing symmetric transverse sets in symcub is to formalize the parallel product with synchronization for process algebra using the associated coskeleton functor (symcub, , Theorem 4.1.8). Indeed, it is impossible to use the coskeleton functor associated with the category of precubical sets because of its pathological behavior (see (ccsprecub, , Proposition 3.15) and (symcub, , Definition 3.1.3)). However, precubical sets still remain sufficient to model this parallel product by tweaking the coskeleton functor of this category (see (ccsprecub, , Section 3.3)).
Symmetric transverse sets share with precubical sets similar metric and homotopical properties by DirectedDegeneracy . Indeed, their geometric realization carries a Lawvere metric structure which enables us to extend Raussen’s notion of (tame) natural $d$-path originally defined for precubical sets (MR2521708, , Definition 2.14) (MR3722069, , Definition 5.3) (MR4070250, , Section 2.9). Moreover, the full subcategory of representable objects of this presheaf category is c-Reedy in the sense of (c-Reedy, , Definition 8.25), like the box category (see Definition 2.3) for the precubical sets (the box category is even direct Reedy in the sense of (ref_model2, , Definition 15.1.2)). This makes possible to compare in (DirectedDegeneracy, , Theorem 7.4) the natural realization of a symmetric transverse set with other realization functors and to generalize homotopical results proved in realization and NaturalRealization for precubical sets.
The technical contribution of this note is threefold. Firstly, we explain why precubical sets and symmetric transverse sets belong to a larger family of presheaf categories on a thick category of cubes (see Definition 2.10). This family of presheaf categories contains also the symmetric precubical sets of MR1988396 and a new category of non-symmetric transverse sets. Symmetric transverse sets are presheaves on a thick category of cubes which turns out to be, for tautological reasons, the greatest one for the inclusion.
@@ 1 Introduction
Secondly, we prove that all results of DirectedDegeneracy are valid for all presheaf categories on a thick category of cubes.
@@
(Section 3 and more specifically Theorem 3.18) All metric and homotopical results of DirectedDegeneracy are valid for the category ${\mathcal{A}}^{op}{{\rm{\mathbf{Set}}}}$ of $\mathcal{A}$-sets when $\mathcal{A}$ is a thick category of cubes.
Thirdly, we obtain a statement which coincides with (a part of) (MR4070250, , Theorem 7.5) when $\mathcal{A}$ is the box category used to define the precubical sets:
@@
(Section 5) The formalization of the parallel composition with synchronisation of symcub using the coskeleton functor of the category of symmetric transverse sets has a category of cube chains which gives the correct space of tame natural $d$-paths up to homotopy.
The formal setting of presheaf category on a thick category of cubes is a first step towards an axiomatization of the notion of tame natural $d$-path. The next step would be to find a way of taking into account the globular version of this notion as it is introduced in Moore3 . This could lead to a general framework unifying all geometric approaches of directed homotopy for concurrency  [^1] The expression “directed homotopy” has several quite distinct meanings. It is the reason why I add “for concurrency” on purpose..
Acknowledgments
I am grateful to the anonymous referee for pointing out the flaw in the proof of Theorem 4.6 (which requires to add Proposition 4.3) and for the suggestions to improve the presentation of the paper.
2 Thick category of cubes
Definition 2.1.
LawvereMetric The small monoidal category $([0,\infty],\geqslant,+,0)$ has for objects the interval $0,\infty]$ and there is a unique arrow $x\to y$ whenever $x\geqslant y$. It is equipped with the monoidal structure induced by addition. A small category enriched over $([0,\infty],\geqslant,+,0)$ is called a Lawvere metric space. The category of Lawvere metric spaces is denoted by $\mathbf{LvMet}$.
Let us now expand Definition 2.1. A Lawvere metric space $(X,d)$ is a set $X$ equipped with a map $d:X\times X\to[0,\infty]$ called a Lawvere metric such that:
@@ 2 Thick category of cubes
A map $f:(X,d)\to(Y,d)$ of Lawvere metric spaces is a function $f:X\to Y$ which is non-expansive, i.e. $\forall(x,y)\in X\times X,d(f(x),f(y))\leqslant d(x,y)$.
Notation 2.2.
@@
Let $0]=\{()\}$ and $n]=\{0<1\}^{n}$ for $n\geqslant 1$ equipped with the product order. Let $0_{n}=(0,\dots,0)$ ($n$ times) and $1_{n}=(1,\dots,1)$ ($n$ times) with $n\geqslant 0$. By convention, one has $\{0<1\}^{0}=[0]=\{()\}$. In the sequel, for all $n\geqslant 1$, both the sets $[n]$ and $0,1]^{n}$ are equipped with the product order. By convention, $0,1]^{0}$ is a singleton.
@@
Let $x=(x_{1},\dots,x_{n})$ and $x^{\prime}=(x^{\prime}_{1},\dots,x^{\prime}_{n})$ be two elements of $0,1]^{n}$ with $n\geqslant 1$. Let $\overrightarrow{d}\!_{1}:[0,1]^{n}\times[0,1]^{n}\to[0,\infty]$ be the function defined by
@@
Let $n\geqslant 0$. The function $\overrightarrow{d}\!_{1}:[0,1]^{n}\times[0,1]^{n}\to[0,\infty]$ is a Lawvere metric by (DirectedDegeneracy, , Proposition 1.5). It restricts to a Lawvere metric on $\{0<1\}^{n}$.
Definition 2.4.
(symcub, , Definition 2.1.5) A map $f:[m]\rightarrow[n]$ of $\rm{\mathbf{PoSet}}$$^{+}$} is cotransverse if
@@
Denote by $\widehat{\square}_{S}$ the subcategory of $\rm{\mathbf{PoSet}}$$^{+}$} consisting of the cotransverse maps.
A cotransverse degeneracy map is a cotransverse map $n]\to[n]$ for $n\geqslant 2$ which is not one-to-one. Proposition 2.5 and Proposition 2.6 are important for the sequel.
Proposition 2.5.
(symcub, , Proposition 3.1.14) Let $0\leqslant m\leqslant n$. Every cotransverse map $f:[m]\rightarrow[n]$ factors uniquely as a composite $m]\stackrel{{\scriptstyle\psi}}{{\longrightarrow}}[m]\stackrel{{\scriptstyle\phi}}{{\longrightarrow}}[n]$ with $\phi\in\square$ and $\psi$ cotransverse.
Proposition 2.6.
Let $f:[m]\to[n]$ be a cotransverse map. Let $\delta:[n]\to[p]$ be a map of $\square$. Suppose that $\delta{f}\in\square$. Then $f\in\square$.
@@
There exists a function $s:[p]\to[n]$ obtained by removing some coordinates such that $s\delta=\operatorname{Id}_{[n]}$. We deduce that $f=s(\delta{f})$. From $\delta{f}\in\square$, we then deduce that $f\in\square$. ∎
Let $\sigma_{i}:[n]\rightarrow[n]$ be the function defined for $1\leqslant i\leqslant n-1$ and $n\geqslant 2$ by
@@
These maps are called the symmetry maps MR1988396 . The symmetry maps are clearly cotransverse. This is the reason of the $S$ in the notation $\widehat{\square}_{S}$.
Notation 2.7.
@@
Definition 2.8.
(symcub, , Definition 2.1.7 and Definition 2.1.12) A category of cubes is a small category $\mathcal{A}$ satisfying the inclusions
@@
The category $\square$, $\square_{S}$ and $\widehat{\square}_{S}$ are examples of categories of cubes. The $\square$-sets are the precubical sets Brown_cube . The $\square_{S}$-sets are the symmetric precubical sets (see MR1988396 ). The $\widehat{\square}_{S}$-sets are the symmetric transverse sets introduced in DirectedDegeneracy .
@@
By (symcub, , Proposition 2.1.15), the identity map $\operatorname{Id}_{\mathcal{L}_{\mathcal{A}}(K)}$ induces for all precubical sets $K$ a natural inclusion of precubical sets $K\subset\omega_{\mathcal{A}}\mathcal{L}_{\mathcal{A}}(K)$. Let $f,g:K\to L$ be two maps of precubical sets such that $\mathcal{L}_{\mathcal{A}}(f)=\mathcal{L}_{\mathcal{A}}(g)$. Then $\omega_{\mathcal{A}}\mathcal{L}_{\mathcal{A}}(f)=\omega_{\mathcal{A}}\mathcal{L}_{\mathcal{A}}(g)$. Thus
@@
Let $K$ be an $\mathcal{A}$-set. The set $K([n])$ is denoted by $K_{n}$. The vertex of $x(0_{n})\in K_{0}$ is called the initial state of the $n$-cube $c$ and the vertex $x(1_{n})\in K_{0}$ is called the final state of the $n$-cube $c$. For any map $k:[m]\rightarrow[n]$ of $\mathcal{A}$ and any $\mathcal{A}$-set $K$, denote by $k^{*}:K_{n}\rightarrow K_{m}$ the function induced by $k$. Let $p\geqslant 0$. The $p$-cube $\mathcal{A}[p]$ is by definition the presheaf $\mathcal{A}(-,[p])$. For any $\mathcal{A}$-set $K$, an element $x\in K_{n}$ corresponds by the Yoneda lemma to a map of $\mathcal{A}$-sets $x:\mathcal{A}[n]\to K$. For any $\mathcal{A}$-set $K$, the data
@@
$\partial\mathcal{A}[n]=\mathcal{A}[n]_{\leqslant n-1}$
@@
The category of cubes $\widehat{\square}_{S}$ is thick for tautological reasons. The terminology must be understood as follows. A thick category of cubes $\mathcal{A}$ is morally a thick subcategory of the category of cubes $\widehat{\square}_{S}$: it is an analogy with the notion of thick subcategory of a triangulated category. Theorem 2.16 provides other examples of thick categories of cubes.
Proposition 2.11.
@@
[…]
@@
[…]
Since $\mathcal{A}$ is thick, write $f^{\prime}=\delta^{\prime\prime}g$ with $\delta^{\prime\prime}\in\square$. We obtain $\delta{f}=k=\delta^{\prime}f^{\prime}=\delta^{\prime}\delta^{\prime\prime}g$. By uniqueness of the factorization of Proposition 2.5, we obtain $\delta=\delta^{\prime}\delta^{\prime\prime}$ and $f=g$. We have obtained the map of $(k{\downarrow}j^{n}_{\mathcal{A}})$
[…]
@@
The first statement is (symcub, , Proposition 2.1.14). Let $n\geqslant 0$. Since $\mathcal{L}_{\mathcal{A}}$ is colimit-preserving, there is a natural map of $\mathcal{A}$-sets ($\square_{<n}$ and $\mathcal{A}_{<n}$ are the full subcategory of $\square$ and $\mathcal{A}$ respectively containing only $0],\dots,[n-1]$):
@@
The above arrow is an isomorphism by Proposition 2.11 and (MR1712872, , Theorem 1 p. 213). ∎
Proposition 2.13.
@@
$\widehat{\square}=\{\phi:[m]\to[n]\in\widehat{\square}_{S}\mid\forall\delta:[p]\to[m]\in\square,\phi\delta\hbox{ one-to-one }\Rightarrow\phi\delta\in\square\}$
is closed under composition and contains all identities of $\widehat{\square}_{S}$. There are the inclusions $\square\subset\widehat{\square}\subset\widehat{\square}_{S}$. In other terms, the set of maps $\widehat{\square}$ yields a well-defined category of cubes. The only one-to-one functions of $\widehat{\square}$ are the maps of $\square$. In particular, the only bijective map of $\widehat{\square}([n],[n])$ is $\operatorname{Id}_{[n]}$ for all $n\geqslant 0$: $\widehat{\square}$ does not contain any symmetry map.
@@
Let $\phi_{1},\phi_{2}\in\widehat{\square}$ such that $\phi_{1}\phi_{2}$ exists. Let $\delta\in\square$ such that $\phi_{1}\phi_{2}\delta$ exists and is one-to-one. Then $\phi_{2}\delta$ is a one-to-one function. Thus $\phi_{2}\delta\in\square$, $\phi_{2}$ belonging to $\widehat{\square}$. We deduce that $\phi_{1}\phi_{2}\delta=\phi_{1}(\phi_{2}\delta)\in\square$ since $\phi_{1}\in\widehat{\square}$. This means that $\phi_{1}\phi_{2}\in\widehat{\square}$. For all $\phi=\operatorname{Id}$, one has $\phi\delta=\delta\in\square$. Hence $\widehat{\square}$ contains all identity maps. Finally suppose that $f:[m]\to[n]\in\widehat{\square}$ is one-to-one. Then $f\operatorname{Id}_{[m]}$ is one-to-one, which implies that $f\in\square$. ∎
Definition 2.14.
The $^\widehat{\square}$-sets are called transverse sets.
The following maps, introduced in (symcub, , Definition 3.1.11), are examples of cotransverse degeneracy maps. Let $\gamma_{i}:[n]\rightarrow[n]$ be the function defined for $1\leqslant i\leqslant n-1$ and $n\geqslant 2$ by
@@
Notation 2.15.
(symcub, , Theorem 3.1.16) The category of cubes generated by the $\delta_{i}^{\alpha}$, $\sigma_{i}$ and $\gamma_{i}$ is denoted by $\overline{\square}$.
Unlike $\widehat{\square}$, the category of cubes $\overline{\square}$ has a conjectural presentation by generators and relations (symcub, , Proposition 3.1.20 and Conjecture 3.1.21).
Theorem 2.16.
The categories of cubes $\square$, $\square_{S}$, $\widehat{\square}$ and $\overline{\square}$ are thick.
@@
Every map $f:[m]\to[n]$ of $\square$ factors uniquely in $\widehat{\square}_{S}$ as a composite $m]\to[m]\to[n]$ such that the right-hand map $m]\to[n]$ belongs to $\square$. Since $f=f\operatorname{Id}_{[m]}$, we deduce by uniqueness that the left-hand map belongs to $\square$. Hence the category $\square$ is thick.
Every map $f:[m]\to[n]$ of $\square_{S}$ factors uniquely in $\widehat{\square}_{S}$ as a composite $m]\to[m]\to[n]$ such that the right-hand map $m]\to[n]$ belongs to $\square$. Since all maps of $\square_{S}$ are one-to-one, the left-hand map $m]\to[m]$ is one-to-one, hence bijective for cardinality reason. We deduce that the left-hand map is a composite of symmetry maps, which means that it belongs to $\square_{S}$. Thus the category $\square_{S}$ is thick.
Consider a map $f:[m]\to[n]$ of $\widehat{\square}$. It factors in $\widehat{\square}_{S}$ as a composite $m]\to[m]\to[n]$ such that the right-hand map $\delta:[m]\to[n]$ belongs to the box category. Denote by $g:[m]\to[m]$ the left-hand map. Let $\delta^{\prime}:[p]\to[m]\in\square$ such that $g\delta^{\prime}$ exists and is one-to-one. Then $f\delta^{\prime}=\delta(g\delta^{\prime})$ is one-to-one, being the composite of two one-to-one functions. Since $f\in\widehat{\square}$, we deduce that $\delta{g}\delta^{\prime}\in\square$. Using Proposition 2.6, we deduce that $g\delta^{\prime}\in\square$. We have proved that $g\in\widehat{\square}$. This means that the category of cubes $\widehat{\square}$ is thick.
@@
Theorem 2.17.
Let $\mathcal{A}$ be a thick category of cubes which does not contain the symmetry maps, i.e. the only bijective map of $\mathcal{A}([n],[n])$ is $\operatorname{Id}_{[n]}$ for all $n\geqslant 0$. Then $\mathcal{A}\subset\widehat{\square}$. In other terms, $\widehat{\square}$ is the greatest thick category of cubes for the inclusion which does not contain any symmetry map.